OpenStudy (kailanineon):

The perimeter of a rectangle is 90 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 8: 7. Which of the following would be the best equation to use to solve this problem? 8x = 7x = 90 8x + 8x + 7x + 7x= 90 8x/7x = 90 2x + 2x=90

OpenStudy (anonymous):

8x + 8x + 7x + 7x = 90 Because the ratio of the length to width is 8:7, you can represent the length as 8x and the width as 7x. Then take the equation of the perimeter of a rectange: 2(length) + 2(width) = 90 2(8x) + 2(7x) = 90 <----------- substitute 16x + 14x = 90

OpenStudy (kailanineon):

Thanks :) That's what I thought it was

OpenStudy (anonymous):

No problem, glad I could help!

OpenStudy (kailanineon):

I appreciate it!